A) A 3.2 g of unknown fluorocarbon compound is decomposed at 298K and produced

310 mL of fluorine gas (F2) is collected over water at 25 °C. What is the percent Fluorine

by mass in this unknown fluorocarbon compound [The vapor pressure of water at 25 °C

is 23.76 mmHg]


B) When 1 mole of this fluorocarbon compound was completely combusted in a bomb

calorimeter, it liberates 3870 kJ of heat which causes the temperature to increase from

25 °C to 26.5 °C. Compute the heat capacity of the bomb calorimeter?


1
Expert's answer
2021-12-17T10:25:02-0500


PV"=nRT," let 1atm"=760mmHg"

"P_{TOTAL}=P_{F_2}+P_{H_2O}"


Let Barometric pressure"=745.4mmHg"


Pressure of Fluorine "=745.4-23.76\\=721.64mmHg"


"P_{F_2}=721.64×\frac{1\>atm}{760\>mmHg}"


"=0.9495atm"


"n=\frac{PV}{RT}=\frac{0.9495×0.310}{0.08210×298}"

"=0.012031"



Mass of "F_2=38\u00d70.012031"

"=0.4572"


"\%" by Mass "=\frac{0.4592}{3.2}×100\%=14.29\%"


Part B


"\Delta\>H=C\Delta\>\theta"

"3870=C\u00d7(26.5-25)"

"C=2580kJ\/K"